OpenStudy (anonymous):

what is the solution g/2=1/5?

OpenStudy (anonymous):

Multiply both sides by 2

OpenStudy (anonymous):

g/2=1/5 5g=2 g=2/5

OpenStudy (anonymous):

g = 2/5

OpenStudy (anonymous):

ignore the guy above he did it a wired way it's ok but just wired. just one step multiply both sides by 2. 2 * (1/5) = 2/5

OpenStudy (anonymous):

ok what i need to know is what is G=?

OpenStudy (anonymous):

\[g = 2/5\]

OpenStudy (anonymous):

if you don't understand why let me know ill try to help out.

OpenStudy (anonymous):

it cant be 2/5 cuz only answer's im given is 4 or 10 or 0.1 or 0.4! i think it is 4 but not sure

OpenStudy (anonymous):

Do you know how to write fractions in decimals?

OpenStudy (anonymous):

divide 2/5 How many times does 5 go into 2? 2/5 = 2.0/5 hop the decimal 2.0/5 = .4

OpenStudy (anonymous):

thank you that is how i did it,but i was not to sure of my answer.(ecofox)!

OpenStudy (anonymous):

Ok, no problem!

OpenStudy (anonymous):

Which two operations would you use to solve the equation 6x + 9 = 30?

OpenStudy (anonymous):

nevermind i got it